# From graph to Riesz continuity

### v2

@@ 4 Polarizations
Polarizations of a Hilbert space.   A decomposition of a Hilbert space $H$ into the orthogonal sum $H=H^{+}\oplus H^{-}$ of two closed infinite-dimensional subspaces is called a polarization of $H$. It is determined by the subspace $H^{+}$ or equivalently by the orthogonal projection $p\in\mathcal{P}^{\star}(H)$ onto $H^{+}$; we will identify polarizations with the corresponding subspaces/projections.
@@
There are two natural equivalence relations for polarizations of $H$, which we call “finite” and “compact”. Two polarizations $p$ and $q$ of $H$ are said to be compactly compatible if $p-q$ is a compact operator; they are said to be finitely compatible if $p-q$ is a finite rank operator (that is, the corresponding subspaces are commensurable).
@@
The compact restricted Grassmanian corresponding to the polarization given by a projection $p$ is defined as the space of all polarizations compatible with $p$:
$\operatorname{Gr}_{K}=\operatorname{Gr}_{K}(H;p)=\left\{q\in\mathcal{P}^{\star}(H)\,|\,q-p\in\mathcal{K}(H)\right\}.$
@@
It is equipped with the norm topology induced from the natural inclusion $\operatorname{Gr}_{K}(H;p)\hookrightarrow\mathcal{P}(H)$, or equivalently the inclusion $\operatorname{Gr}_{K}(H;p)\hookrightarrow\mathcal{K}(H)$, $q\mapsto q-p$. The finite restricted Grassmanian $\operatorname{Gr}_{f}(H;p)$ is defined in a similar manner; it is a subspace of $\operatorname{Gr}_{K}(H;p)$. The restricted Grassmanian depends only on the corresponding equivalence class of the projection $p$.
@@
In general, a subbundle $\mathcal{E}$ of $\mathcal{H}$ is given by a family of subspaces $\mathcal{E}_{x}\subset\mathcal{H}_{x}$ which is continuous in an appropriate sense. We will use the following notion of continuity: we say that $\mathcal{E}$ is a subbundle of $\mathcal{H}$ if $\mathcal{E}$ is the range of a bundle projection, that is, a continuous bundle map $p\colon\mathcal{H}\to\mathcal{H}$ such that $p_{x}\in\mathcal{P}(\mathcal{H}_{x})$ for every $x\in X$. In this case, $\mathcal{H}\ominus\mathcal{E}=\operatorname{Ran}(1-p)$ is also a subbundle. If both $\mathcal{E}$ and $\mathcal{H}\ominus\mathcal{E}$ are subbundles of $\mathcal{H}$ in the sense of [DD], then they are subbundles in our sense. See [DD], Proposition 16.
@@
Equivalently, for every $x\in X$, some (and then every) local trivialization of $\mathcal{H}$ over a neighborhood $X^{\prime}$ of $x$ takes $p$ to a continuous map $X^{\prime}\to{{}^{s}{\mathcal{P}}}(H)$, where ${{}^{s}{\mathcal{P}}}(H)$ denotes the space $\mathcal{P}(H)$ equipped with the strong operator topology. In particular, subbundles of the trivial Hilbert bundle $H_{X}=X\times H$ are in one-to-one correspondence with continuous maps $X\to{{}^{s}{\mathcal{P}}}(H)$.
@@
Two subbundles of $\mathcal{H}$ given by bundle projections $p$ and $q$ are said to be (compactly) compatible if $p-q$ is a norm continuous family of compact operators. They are said to be finitely compatible if additionally all operators $p_{x}-q_{x}$ have finite rank.
Polarizations of a Hilbert bundle.   We need only compact polarizations in this paper, so we will mostly use the term “polarization” without the adjective “compact”. The changes required to obtain the finite version are obvious.
@@
Locally, a polarization of $\mathcal{H}$ is determined by a decomposition of $\mathcal{H}$, over some open subset $X^{\prime}\subset X$, into the orthogonal sum $\mathcal{H}=\mathcal{H}^{+}\oplus\mathcal{H}^{-}$ of two subbundles of infinite rank. Equivalently, it is determined by the subbundle $\mathcal{H}^{+}$ of $\left.\mathcal{H}\right|_{X^{\prime}}$, or by the bundle projection with the range $\mathcal{H}^{+}$. By definition, two compatible subbundles of $\mathcal{H}$ over $X^{\prime}$ determine the same polarization.
Globally, a polarization of $\mathcal{H}$ is determined by an open covering $(X_{i})$ of $X$ and subbundles $\mathcal{H}^{+}_{i}$ of $\left.\mathcal{H}\right|_{X_{i}}$ such that $\mathcal{H}^{+}_{i}$ has infinite rank and corank and is compatible with $\mathcal{H}^{+}_{j}$ over $X_{i}\cap X_{j}$ for every $i,j$. Following Ivanov [$I_{2}$], we call such a collection $(X_{i},\mathcal{H}^{+}_{i})$ a polarization atlas. Two polarization atlases are said to be equivalent if their union is also a polarization atlas. A polarization of $\mathcal{H}$ is an equivalence class of polarization atlases.
@@
A polarization $\Pi$ of $\mathcal{H}$ determines the restricted Grassmanian bundle $\operatorname{Gr}_{K}(\mathcal{H};\Pi)$ in an obvious way. This is one of the motivations behind the notion of a polarization.
@@
Two polarizations of $\mathcal{H}$ are said to be homotopic if they are restrictions to $X\times\left\{0\right\}$ and $X\times\left\{1\right\}$ of some polarization of the Hilbert bundle ×$\mathcal{H}\times[0,1]\to X\times[0,1]$.
@@
Adapted trivializations.   Following [$I_{2}$], we say that a trivialization $\mathcal{H}\to X\times H$ is adapted to a subbundle $\mathcal{E}$ of $\mathcal{H}$ if it takes the family of subspaces $\mathcal{E}_{x}\subset\mathcal{H}_{x}$ to a continuous map $X\to\operatorname{Gr}(H)$. If a trivialization of $\mathcal{H}$ is adapted to $\mathcal{E}$, then it is adapted to every subbundle compatible with $\mathcal{E}$.
@@
Let $\Pi$ be a polarization given by an atlas $(X_{i},\mathcal{E}_{i})$. We say that a trivialization of $\mathcal{H}$ is adapted to $\Pi$ if its restrictions to $X_{i}$ are adapted to $\mathcal{E}_{i}$. The notion of an adapted local trivialization is defined in an obvious way. We call a polarization $\Pi$ of $\mathcal{H}$ tame if it admits local adapted trivializations (that is, for every point $x\in X$ there is a local trivialization of $\mathcal{H}$ adapted to $\Pi$ over some neighborhood of $x$).
The notion of a polarization introduced by Ivanov in the appendix to the third version of his preprint [$I_{2}$] is a particular case of a tame finite polarization in our terminology.

### v3

@@ 5 Polarizations and proof of Theorem 4
Polarizations of a Hilbert space.  A decomposition of a Hilbert space $H$ into the orthogonal sum $H=H^{+}\oplus H^{-}$ of two closed infinite-dimensional subspaces is called a polarization of $H$. It is determined by the subspace $H^{+}\in\mathcal{G}r^{\star}(H)$ or equivalently by the orthogonal projection onto $H^{+}$. We denote by $\mathcal{P}^{\star}(H)\cong\mathcal{G}r^{\star}(H)$ the subspace of $\mathcal{P}(H)$ consisting of projections of infinite rank and corank and will identify polarizations with the corresponding subspaces/projections.
@@
There are two natural equivalence relations for polarizations of $H$, which we call “finite” and “compact”. Two polarizations $p$ and $q$ of $H$ are said to be compactly compatible if pqp-q is a compact operator; they are said to be finitely compatible if pqp-q is a finite rank operator (that is, the corresponding subspaces are commensurable).
@@
The compact restricted Grassmanian corresponding to a polarization $p$ is defined as the space of all polarizations compatible with $p$:
$\mathcal{G}r_{K}(H;p)=\left\{q\in(H)\mid q-p\in\mathcal{K}(H)\right\}.$
@@
It is equipped with the norm topology induced by the natural inclusion $\mathcal{G}r_{K}(H;p)\hookrightarrow\mathcal{G}r(H)\cong\mathcal{P}(H)$, or equivalently by the inclusion $\mathcal{G}r_{K}(H;p)\hookrightarrow\mathcal{K}(H)$, $q\mapsto q-p$. The finite restricted Grassmanian $\mathcal{G}r_{f}(H;p)$ is defined in a similar manner; it is a subspace of $\mathcal{G}r_{K}(H;p)$. The restricted Grassmanian depends only on the corresponding equivalence class of the projection $p$.
@@
In general, a subbundle $\mathcal{E}$ of $\mathcal{H}$ is given by a family of closed subspaces $\mathcal{E}_{x}\subset\mathcal{H}_{x}$ which is continuous in an appropriate sense. We will use the following notion of continuity: we say that $\mathcal{E}$ is a subbundle of $\mathcal{H}$ if $\mathcal{E}$ is the range of a bundle projection, that is, a continuous bundle map $p\colon\mathcal{H}\to\mathcal{H}$ such that $p_{x}\in\mathcal{P}(\mathcal{H}_{x})$ for every $x\in X$. In this case $\mathcal{H}\ominus\mathcal{E}=\operatorname{Ran}(1-p)$ is also a subbundle.
@@
Equivalently, for every $x\in X$, some (and then every) local trivialization of $\mathcal{H}$ over a neighborhood $Y$ of $x$ takes $p$ to a continuous map $Y\to{{}^{s}{\mathcal{P}}}(H)$, where ${{}^{s}{\mathcal{P}}}(H)$ denotes the space of projections equipped with the strong operator topology. In particular, subbundles of the trivial Hilbert bundle $H_{X}=X\times H$ are in one-to-one correspondence with continuous maps $X\to{{}^{s}{\mathcal{P}}}(H)$.
A family $\mathcal{E}$ of closed subspaces $\mathcal{E}_{x}\subset\mathcal{H}_{x}$ is a subbundle of $\mathcal{H}$ in our sense if and only if both $(\mathcal{E}_{x})$ and $(\mathcal{H}_{x}\ominus\mathcal{E}_{x})$ are continuous subfields of the field $(\mathcal{H}_{x})$ of Hilbert spaces in the sense of Dixmier and Douady [DD]. See [DD], Proposition 16.
@@
We say that two subbundles of $\mathcal{H}$ given by bundle projections $p$ and $q$ are (compactly) compatible if pqp-q is a norm continuous family of compact operators; they are finitely compatible if additionally all operators $p_{x}-q_{x}$ have finite rank.
Polarizations of a Hilbert bundle.  A natural forgetful functor from finite to compact polarizations preserves the notion of an adapted trivialization. Therefore, the “finite” analogues of Theorems 4 and 5 follow immediately from these theorems. Because of that, we will deal only with compact polarizations in the rest of the paper and will mostly omit the adjective “compact” before “polarization”. The changes required to obtain the finite version of our statements are obvious.
@@
Locally, a polarization of $\mathcal{H}$ is determined by a decomposition of $\mathcal{H}$, over some open subset $Y\subset X$, into the orthogonal sum $\mathcal{H}=\mathcal{H}^{+}\oplus\mathcal{H}^{-}$ of two subbundles of infinite rank. Equivalently, it is determined by the subbundle $\mathcal{H}^{+}$ of $\left.\mathcal{H}\right|_{Y}$ or by the bundle projection with the range $\mathcal{H}^{+}$. By the definition, two compatible subbundles of $\mathcal{H}$ over $Y$ determine the same local polarization.
[…]
@@
A polarization $\Pi$ of $\mathcal{H}$ determines the restricted Grassmanian bundle $\mathcal{G}r_{K}(\mathcal{H};\Pi)$ in an obvious way. This is one of the motivations behind the notion of a polarization.
@@
Two polarizations of $\mathcal{H}$ are said to be homotopic if they are restrictions to $X\times\left\{0\right\}$ and $X\times\left\{1\right\}$ of some polarization of the Hilbert bundle $\mathcal{H}\times[0,1]\to X\times[0,1]$.
@@
Adapted trivializations.  A trivialization $\mathcal{H}\to X\times H$ is said to be adapted to a subbundle $\mathcal{E}$ of $\mathcal{H}$ if it takes the family of subspaces $\mathcal{E}_{x}\subset\mathcal{H}_{x}$ to a continuous map $X\to\mathcal{G}r(H)$, or equivalently takes the corresponding family of projections to a norm continuous map $X\to\mathcal{P}(H)$. If a trivialization of $\mathcal{H}$ is adapted to $\mathcal{E}$, then it is adapted to every subbundle compatible with $\mathcal{E}$, since the difference pqp-q of corresponding bundle projections is norm continuous in every trivialization.
@@
Let $\Pi$ be a polarization given by an atlas $(X_{i},\mathcal{E}_{i})$. We say that a trivialization of $\mathcal{H}$ is adapted to $\Pi$ if its restriction to $X_{i}$ is adapted to $\mathcal{E}_{i}$ for every $i$. The notion of an adapted local trivialization is defined in an obvious way.
We call a subbundle $\mathcal{E}$, resp. a polarization $\Pi$ of $\mathcal{H}$ tame if it admits local adapted trivializations (that is, for every point $x\in X$ there is a local trivialization of $\mathcal{H}$ adapted to $\mathcal{E}$, resp. $\Pi$ over some neighborhood of $x$).
Trivializations of $H_{X}$ adapted to $E_{X}$.  As a first step to the proof of Theorem 4, we describe trivializations of the trivial Hilbert bundle $H_{X}=X\times H$ adapted to its trivial subbundle $E_{X}=X\times E$, where $E$ is a fixed closed subspace of $H$.
Trivializations of $H_{X}$ are given by continuous maps $f\colon X\to{{}^{s}{\mathcal{U}}}(H)$. Such a trivialization is adapted to $E_{X}$ if and only if the map $X\to\mathcal{G}r(H)$ taking $x$ to $f(x)E$ is continuous. Therefore, trivializations of $H_{X}$ adapted to $E_{X}$ are in one-to-one correspondence with continuous maps $X\to T_{E}$, where the space $T_{E}$ is obtained from ${{}^{s}{\mathcal{U}}}(H)$ by refining its topology so that the map
(5.1) $\iota_{E}\colon T_{E}\to\mathcal{G}r(H),\quad u\mapsto uE$
becomes continuous. In other words, $T_{E}$ has the topology induced by the injection
$T_{E}\hookrightarrow{{}^{s}{\mathcal{U}}}(H)\times\mathcal{G}r(H),\quad u\mapsto(u,uE).$
Two maps $f,g\colon X\to T_{E}$ are homotopic if and only if the corresponding trivializations are homotopic in the class of trivializations adapted to $E_{X}$.
Lemma 5.1.
If $E\in\mathcal{G}r^{\star}(H)$, then $T_{E}$ is contractible.
Proof. For $E\in\mathcal{G}r^{\star}$ the map (5.1) takes $T_{E}$ to $\mathcal{G}r^{\star}$. We will show that $\iota_{E}\colon T_{E}\to\mathcal{G}r^{\star}$ is a trivial bundle and construct its trivialization.
The base space $\mathcal{G}r^{\star}$ has a distinguished point $E\in\mathcal{G}r^{\star}$. The fiber of $\iota_{E}$ over this point is the subspace of ${{}^{s}{\mathcal{U}}}(H)$ consisting of unitary operators $u\colon H\to H$ preserving $E\subset H$. Denote this subspace by ${{}^{s}{\mathcal{U}}}_{E}$. Let $g\colon\mathcal{G}r^{\star}\to\mathcal{U}(H)$ be a continuous map such that $L=g_{L}E$ for all $L\in\mathcal{G}r^{\star}$, as in (2.1). Consider the map
$\varphi\colon{{}^{s}{\mathcal{U}}}_{E}\times\mathcal{G}r^{\star}\;\to\;T_{E},\quad\varphi(v,L)=g_{L}\cdot v.$
The composition $\iota_{E}\circ\varphi\colon{{}^{s}{\mathcal{U}}}_{E}\times\mathcal{G}r^{\star}\to\mathcal{G}r^{\star}$ is the projection of the product on the second factor, as can be seen from
(5.2) $(g_{L}v)E=g_{L}E=L\;\;\text{for every}\;\;v\in{{}^{s}{\mathcal{U}}}_{E}.$
In particular, $\iota_{E}\circ\varphi$ is continuous. The composition of $\varphi$ with the identity map $T_{E}\to{{}^{s}{\mathcal{U}}}$ is given by $(v,L)\mapsto g_{L}v$ and is also continuous. Therefore, $\varphi$ itself is continuous.
$\psi\colon T_{E}\to{{}^{s}{\mathcal{U}}}_{E}\times\mathcal{G}r^{\star},\quad\psi(u)=\left(g_{uE}^{-1}\cdot u,\,uE\right).$
By the construction, $\psi\circ\varphi(v,L)=(v,L)$ for every $(v,L)\in{{}^{s}{\mathcal{U}}}_{E}\times\mathcal{G}r^{\star}$. Conversely,
$\varphi\circ\psi(u)=\varphi\left(g_{uE}^{-1}\cdot u,\,uE\right)=g_{uE}\cdot g_{uE}^{-1}\cdot u=u\;\;\text{for every}\;\;u\in T_{E}.$
The domain of $\varphi$ is the product of two factors. The second factor $\mathcal{G}r^{\star}$ is contractible, see [AS, proof of Lemma 3.6]. The first factor ${{}^{s}{\mathcal{U}}}_{E}={{}^{s}{\mathcal{U}}}(E)\times{{}^{s}{\mathcal{U}}}(E^{\bot})$ is the product of two copies of $^{s}{\mathcal{U}}$}. By [DD, Lemma 3], $^{s}{\mathcal{U}}$} is contractible, so ${{}^{s}{\mathcal{U}}}_{E}$ is also contractible. Therefore, the product ${{}^{s}{\mathcal{U}}}_{E}\times\mathcal{G}r^{\star}\cong T_{E}$ is contractible. $\square$
Trivializations of $\mathcal{H}$.  Let $\mathcal{H}$ be a locally trivial Hilbert bundle and $H_{X}$ be the trivial Hilbert bundle with the fiber $H$ over a topological space $X$. Let $\mathcal{U}(\mathcal{H},H_{X})$ be the fiber bundle associated with $\mathcal{H}$ whose fiber over $x$ is ${{}^{s}{\mathcal{U}}}(\mathcal{H}_{x},H)$. This is a locally trivial principal bundle over $X$ with the structure group ${{}^{s}{\mathcal{U}}}(H)$ acting on fibers by the left multiplication. Sections of $\mathcal{U}(\mathcal{H},H_{X})$ are Hilbert bundle isomorphisms $\mathcal{H}\to H_{X}$, that is, trivializations of $\mathcal{H}$. A trivialization $\mathcal{H}\to H_{X}$ takes $\mathcal{U}(\mathcal{H},H_{X})$ to the trivial bundle $\mathcal{U}(H_{X},H_{X})=X\times{{}^{s}{\mathcal{U}}}(H)$.
Trivializations adapted to $\mathcal{E}$.  Let $\mathcal{E}$ be a subbundle of $\mathcal{H}$. Consider the space $\mathcal{T}=\mathcal{T}_{\mathcal{E}}$ which coincides with $\mathcal{U}(\mathcal{H},H_{X})$ as a set, but has finer topology determined by $\mathcal{E}$. Namely, we equip $\mathcal{T}$ with the weakest topology such that both the identity map $\iota\colon\mathcal{T}\to\mathcal{U}(\mathcal{H},H_{X})$ and the map
(5.3) $\iota_{\mathcal{E}}\colon\mathcal{T}\to\mathcal{G}r(H)\cong\mathcal{P}(H),\quad\iota_{\mathcal{E}}(x,u)\ =u(\mathcal{E}_{x})\;\;\text{for}\;\;x\in X\;\;\text{and}\;\;u\in\mathcal{U}(\mathcal{H}_{x},H)$
are continuous. Then the composition $\mathcal{T}\to\mathcal{U}(\mathcal{H},H_{X})\to X$ is also continuous, and we consider $\mathcal{T}$ as a space over $X$. By the construction, sections of $\mathcal{T}\to X$ are in one-to-one correspondence with trivializations of $\mathcal{H}$ adapted to $\mathcal{E}$.
Lemma 5.2.
If $\mathcal{E}$ and $\mathcal{F}$ are compatible subbundles of $\mathcal{H}$, then $\iota_{\mathcal{F}}$ and $\iota_{\mathcal{E}}$ induce the same refinement of the original topology on $\mathcal{U}(\mathcal{H},H_{X})$. In other words, $\mathcal{T}_{\mathcal{E}}$ depends only on the equivalence class of compatible subbundles $\mathcal{E}$.
Proof. Let $p_{x}$ and $q_{x}$ be projections onto $\mathcal{E}_{x}$ and $\mathcal{F}_{x}$ respectively. By the definition of compatible subbundles, the difference $k_{x}=q_{x}-p_{x}$ is a norm continuous family of compact operators, so $\iota_{\mathcal{F}}$ and $\iota_{\mathcal{E}}$ considered as maps from $\mathcal{U}(\mathcal{H},H_{X})$ to $\mathcal{P}(H)$ differ by a continuous map
$\mathcal{U}(\mathcal{H},H_{X})\to\mathcal{K}(H),\quad(x,u)\mapsto\iota_{\mathcal{F}}(x,u)-\iota_{\mathcal{E}}(x,u)=uq_{x}u^{*}-up_{x}u^{*}=uk_{x}u^{*}.$
Hence $\iota_{\mathcal{F}}$ and $\iota_{\mathcal{E}}$ induce the same refinement of the original topology on $\mathcal{U}(\mathcal{H},H_{X})$. $\square$
Lemma 5.3.
If $\mathcal{E}$ is tame, then $\mathcal{T}_{\mathcal{E}}\to X$ is a locally trivial bundle, whose fiber $T=T_{E}$ is the space defined before Lemma 5.1.
Proof. Let $x_{0}\in X$. Since $\mathcal{E}$ is tame, there is a trivialization $\tau^{\prime}$ of $\mathcal{H}$ adapted to $\mathcal{E}$ over a neighborhood $X^{\prime}$ of $x_{0}$. This trivialization takes $\mathcal{E}$ to the subbundle $\mathcal{E}^{\prime}$ given by a continuous map $X^{\prime}\to\mathcal{G}r(H)$. Replacing $X^{\prime}$ by a smaller neighborhood if necessary, we can find a map $g\colon X^{\prime}\to\mathcal{U}(H)$ such that $g_{x}\mathcal{E}^{\prime}_{x}=E$, where $E=\mathcal{E}^{\prime}_{x_{0}}$. The composition of $\tau^{\prime}$ with the automorphism of the trivial Hilbert bundle $H_{X^{\prime}}$ provided by $g$ determines the new trivialization $\tau$ which takes $\mathcal{E}_{x}$ to the subspace $E$ of $H$ independent of $x$ and thus takes $\mathcal{E}$ to the constant subbundle $E_{X^{\prime}}$ of $H_{X^{\prime}}$.
Such a trivialization $\tau$ provides a bundle isomorphism $\left.\mathcal{T}_{\mathcal{E}}\right|_{X^{\prime}}\to\mathcal{T}^{\prime}$, where the bundle $\mathcal{T}^{\prime}$ is build (in the way described above) from the trivial bundle $H_{X^{\prime}}$ and its subbundle $E_{X^{\prime}}$. By the definition, the total space $\mathcal{T}^{\prime}$ is obtained from the product $X^{\prime}\times{{}^{s}{\mathcal{U}}}(H)$ by refining its topology so that the map $\mathcal{T}^{\prime}\to\mathcal{G}r(H)$, $(x,u)\mapsto uE$ becomes continuous. But this map does not actually depend on $x$, so $\mathcal{T}^{\prime}$ coincides with the trivial bundle $X^{\prime}\times T_{E}\to X^{\prime}$. $\square$
Trivializations adapted to $\Pi$.  Let $\Pi$ be a polarization of $\mathcal{H}$ given by an atlas $(X_{i},\mathcal{E}_{i})$, $i\in I$. Consider the space $\mathcal{T}_{\Pi}$ which coincides with $\mathcal{U}(\mathcal{H},H_{X})$ as a set and is equipped with the weakest topology such that the identity map
(5.4) $\left.\mathcal{T}_{\Pi}\right|_{X_{i}}\to\mathcal{T}_{\mathcal{E}_{i}}$
is continuous for every $i\in I$. By Lemma 5.2 the restrictions of $\mathcal{T}_{\mathcal{E}_{i}}$ and $\mathcal{T}_{\mathcal{E}_{j}}$ to $X_{i}\cap X_{j}$ coincide, so (5.4) is actually a homeomorphism. Moreover, this definition does not depend on the choice of an atlas of $\Pi$.
The composition $\mathcal{T}_{\Pi}\to\mathcal{U}(\mathcal{H},H_{X})\to X$ is continuous, and we consider $\mathcal{T}_{\Pi}$ as a space over $X$. By the construction, (local or global) sections of $\mathcal{T}_{\Pi}\to X$ are exactly (local or global) trivializations of $\mathcal{H}$ adapted to $\Pi$. In particular, $\mathcal{T}_{\Pi}\to X$ admits local sections if and only if $\Pi$ is tame.
Proof of Theorem 4.  Let $\Pi$ be a tame polarization of a locally trivial Hilbert bundle $\mathcal{H}$ over $X$. Then, by Lemma 5.3, $\mathcal{T}_{\Pi}\to X$ is a locally trivial bundle with the fiber $T_{E}$, $E\in\mathcal{G}r^{\star}(H)$. This fiber is contractible by Lemma 5.1. If additionally $X$ is paracompact, then $\mathcal{T}_{\Pi}\to X$ admits global sections, every two sections are homotopic, and a section over a closed subspace $Y\subset X$ can be extended to a global section [DD, Lemma 4]. Since sections of $\mathcal{T}_{\Pi}$ correspond to adapted trivializations, this proves the theorem. $\square$
6 Proof of Theorem 5
Let $\dim Z$ denote the covering dimension of a paracompact space $Z$. Recall that a closed subset $Y$ of a paracompact space $Z$ is itself paracompact [E, Corollary 5.1.29], normal [E, Theorem 5.1.5], and $\dim Y\leqslant\dim Z$ [E, Theorem 7.1.8].
Finite covering dimension.  Our proof of Theorem 5 is based on Theorem 6.1 below. Theorem 6.1 follows from a theorem of Dixmier and Douady [DD, Theorem 5]. For the convenience of the reader, we include here a proof of Theorem 6.1; it is based on the same ideas as the proof of [DD, Theorem 5], but is stated in terms of operators in a fixed Hilbert space instead of “continuous fields of Hilbert spaces” of Dixmier–Douady.
Let $E$ be a subspace of $H$ of infinite dimension and codimension and $p_{E}\in\mathcal{P}^{\star}(H)$ be the orthogonal projection onto $E$. In terms of projections, the map $\mathcal{U}(H)\to\mathcal{G}r^{\star}(H)$, $u\mapsto uE$ is written as $\mathcal{U}(H)\to\mathcal{P}^{\star}(H)$, $u\mapsto up_{E}u^{-1}$. This map has a global section, as explained in the beginning of Section 2. In contrast with this, the continuous map
${{}^{s}{\mathcal{U}}}(H)\to{{}^{s}{\mathcal{P}}}^{\star}(H)$
given by the same formula $u\mapsto up_{E}u^{-1}$ has no sections, even locally. However, the following result holds.